OpenStudy (ny,ny):

What sequence of stars matches the rule s=1+(2n+1) where n represents the term number and s represents the number of stars?

OpenStudy (ny,ny):

answer choice a: term 1: 1 star term 2: 2 stars term 3: 3 stars

OpenStudy (ny,ny):

answer choice b: term 1: 3 stars term 2: 5 stars term 3: 7 stars

OpenStudy (ny,ny):

answer choice c: term 1: 4 stars term 2: 8 stars term 3: 11 stars

OpenStudy (ny,ny):

answer choice d: term 1: 4 stars term 2: 6 stars term 3: 8 stars

OpenStudy (ny,ny):

i do not understand plz help

OpenStudy (anonymous):

Here is what this means... n is the term.. so you would put in 1 for n and find s When n = 1, s = 1 + (2x1 + 1) you do parentheses first so when n = 1 s = 4 When n = 2 s = 1 + (2x2+1) so when n = 2, s = 6 When n = 3, s = 1 + (2x3 + 1) so when n = 3, s = 8

OpenStudy (ny,ny):

so answer choice D?

OpenStudy (anonymous):

Yes.. do you understand why?

OpenStudy (ny,ny):

yes ty very much